OpenStudy (anonymous):

what are the critical values of f'(x)= x^2+9/x^2

OpenStudy (anonymous):

critical points are those where the derivative vanishes

OpenStudy (anonymous):

f'(x)=0

OpenStudy (anonymous):

or the derivative does not exist

OpenStudy (anonymous):

ok f we see denominator x^2 is 0,but in the up x^2+9 if we see c^2+9=0 then c^2=-9 which is 3i i dont know the solution here

OpenStudy (anonymous):

so it means that c=0 is one critical value,coz derivative is not defined there

OpenStudy (anonymous):

but do mind one thing,function should be defined at the critical value

OpenStudy (anonymous):

so what do you think the solution here , is have any critical value ?

OpenStudy (anonymous):

what function is actually given in this prob,u posted the derivative

OpenStudy (anonymous):

f(x)=9/x+x is the funtion